# Mode analysis of Nambu-Goldstone modes in U(1) charged first-order relativistic viscous hydrodynamics

### v2

@@ 2.6 Hydrodynamical on-shell transformation
To derive Eqs. (102) and (103), we have exhausted the freedom of hydrodynamical frame transformations. The transport coefficients with a tilde are hydrodynamical frame invariants, so we cannot further eliminate them by the hydrodynamical frame transformation. However, when evaluated on the lowest order hydrodynamical solutions, one can further simplify the coefficients Kovtun:2012rj . The lowest-order hydrodynamical equations
$\nabla_{\mu}(nu^{\mu})$ $=0,$ (111)
$u_{\nu}\nabla_{\rho}(\rho u^{\rho}u^{\nu}+P\gamma^{\rho\nu})$ $=0,$ (112)
$\gamma_{\mu\nu}\nabla_{\rho}(\rho u^{\rho}u^{\nu}+P\gamma^{\rho\nu})+\gamma_{\mu\nu}nu_{\rho}F^{\rho\nu}$ $=0,$ (113)
are recast into
$u^{\rho}\nabla_{\rho}\beta$ $=\frac{n\rho_{\mu}-n_{\mu}(\rho+P)}{n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu}}\nabla_{\rho}u^{\rho},$ (114)
$u^{\rho}\nabla_{\rho}(\beta\mu)$ $=\frac{-\beta n\rho_{\beta}+\mu n\rho_{\mu}+\beta n_{\beta}(\rho+P)-\mu n_{\mu}(\rho+P)}{\beta(n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu})}\nabla_{\rho}u^{\rho},$ (115)
$u^{\rho}\nabla_{\rho}u_{\mu}$ $=-\frac{P_{\beta}-\frac{\mu}{\beta}P_{\mu}}{\rho+P}\gamma_{\mu}{}^{\nu}\nabla_{\nu}\beta-\frac{P_{\mu}}{\beta(\rho+P)}\gamma_{\mu}{}^{\nu}\nabla_{\nu}(\beta\mu)-\frac{n}{\rho+P}u^{\rho}F_{\rho\mu}.$ (116)
@@
$P_{(1)}$ $={\bm{\zeta}}\nabla_{\rho}u^{\rho},$ (117)
$j^{\mu}_{(1)}$ $={\bm{\tau}}\gamma^{\mu\nu}\nabla_{\nu}\beta-{\bm{\chi}}_{1}\beta\gamma^{\mu\nu}\nabla_{\nu}(\beta\mu)-{\bm{\chi}}_{2}\beta^{2}\gamma^{\mu\nu}u^{\rho}F_{\rho\nu},$ (118)
${\bm{\zeta}}$ $\equiv\tilde{\zeta}+\left(\frac{\partial P}{\partial\rho}\right)_{n}\tilde{\lambda}+\left(\frac{\partial P}{\partial n}\right)_{\rho}\tilde{\nu},$ (119)
${\bm{\tau}}$ $\equiv\tilde{\tau}_{1}-\tilde{\tau}_{2},$ (120)
${\bm{\chi}}_{1}$ $\equiv\tilde{\chi}_{1}-\frac{n}{\beta(\rho+P)}\tilde{\tau}_{2},$ (121)
${\bm{\chi}}_{2}$ $\equiv\tilde{\chi}_{2}-\frac{n}{\beta(\rho+P)}\tilde{\tau}_{2},$ (122)
where we used Eqs. (285) and (287). The $\bm{\zeta}$, $\bm{\chi}_{i}$ and $\bm{\tau}_{i}$ are hydrodynamical frame invariants by definition.
When imposing the KMS condition in the original hydrodynamical frame as presented in App. A.2, one can further constrain the coefficients as
${\bm{\tau}}=0,~{}{\bm{\chi}}_{1}={\bm{\chi}}_{2}\equiv{\bm{\chi}}.$ (123)
@@
$x$ $=1,~{}y=-\left(\frac{\partial P}{\partial\rho}\right)_{n},~{}z=-\left(\frac{\partial P}{\partial n}\right)_{\rho},$ (125)
in Eq. (73) and
$x=-\frac{n}{\beta(\rho+P)},~{}y=1,$ (126)
in Eq. (72). Now, we arrive at the so-called Landau frame, where the first-order transport coefficients are characterized by three independent hydrodynamic parameters: the shear viscosity $\eta$, bulk viscosity $\zeta$, and charge conductivity $\chi$.
In Ref. Kovtun:2019hdm , the ambiguity of the on-shell condition was emphasized. Specifically, it was shown that one can eliminate the term $\nabla_{\mu}u^{\mu}$ and express the constitutive relations in terms of either $u^{\mu}\nabla_{\mu}\beta$ or $u^{\mu}\nabla_{\mu}(\beta\mu)$. These correspond to the conditions $\tilde{\lambda}=\tilde{\zeta}=\tilde{\tau}=0$ and $\tilde{\zeta}=\tilde{\nu}=\tilde{\tau}=0$, respectively, thus modifying the hydrodynamic equations of motion. Therefore, on-shell solutions reintroduce ambiguities in the transport coefficients, which we refer to as on-shell transformations.
@@ 2.7 Traditional hydrodynamical frames
The above derivation of the Landau frame differs from the original prescription given by Landau-Lifshitz, which we discuss below. Suppose one imposes the KMS condition after maximally eliminating the coefficients via field redefinition, which corresponds to taking $T_{(1)}^{\mu\nu}u_{\nu}=j_{(1)}^{\mu}u_{\mu}=0$, followed by the imposition of the second law of thermodynamics as outlined in Landau and Lifshitz’s textbook. The first condition then leads to $\epsilon=\lambda_{1}=\nu_{1}=\nu_{2}=\nu_{3}=\nu_{4}=\kappa_{1}=\kappa_{2}=\tau_{3}=\tau_{4}=0$. The second condition yields $\lambda_{2}=\lambda_{1}$, $\nu_{5}=\nu_{2}$, $\kappa_{i}=\kappa$, $\tau_{i}=\tau$, and $\chi_{i}=\chi$. As a result, $T^{\mu\nu}$ and $j^{\mu}$ are expressed in terms of three non-vanishing transport coefficients: $\eta$, $\zeta$, and $\chi$. The noise coefficients, except for $w_{4}=\zeta$, $w_{5}=\eta$, and $y_{2}=\chi$, vanish by the KMS condition.
One may identify $(\zeta,\chi)$ with $({\bm{\zeta}},\bm{\chi})$. However, we point out that five independent noise coefficients remain in the method in the previous subsection, while only three exist in the prescription of Landau-Lifshitz. This inconsistency suggests that on-shell hydrodynamics, including the noise field, must be solved. We will revisit this point in Sec. 5, where we will show that the noise coefficients can also be reduced.
Similarly, the Eckart frame is realized by imposing $T_{(1)}^{\mu\nu}u_{\nu}u_{\mu}=j_{(1)}^{\mu}=0$, which corresponds to $\epsilon=\lambda_{1}=\nu_{1}=\nu_{2}=\nu_{3}=\nu_{4}=\tau_{1}=\tau_{2}=\chi_{1}=\chi_{2}=0$. Then, with the KMS condition, the frame invariants are set to $\tilde{\lambda}=\tilde{\nu}=\tilde{\chi}=0$, and $\tilde{\zeta}=\zeta$ and $\tilde{\tau}=-n\kappa/\beta(\rho+P)$. Note that the traditional Landu frame is equivalent to taking the special condition $\tilde{\lambda}=\tilde{\nu}=\tilde{\tau}=0$ and $\tilde{\zeta}=\zeta$ and $\tilde{\chi}=\chi$ in the most general hydrodynamics. Therefore, these frames are not equivalent under hydrodynamical transformations, as these frames have different hydrodynamical frame invariants. Landau and Eckart frames are equivalent up to the hydrodynamical frame transformation followed by the on-shell transformation.
At first glance, the lowest-order hydrodynamics can be solved arbitrarily, making on-shell transformation seem like an unphysical freedom. Then, the freedom stems from a combination of arbitrary frame transformations, and on-shell transformation should be unphysical. However, the on-shell transformation changes the nature of the equations of motion, which has implications for stability and causality analysis Kovtun:2019hdm . Previous works have concluded that traditional first-order hydrodynamics is neither stable nor causal because of this uncertainty. In the next section, we will demonstrate that mode analysis remains unaffected by this freedom.
@@ 2.8 Ambiguity of rest frames
In hydrodynamics, “frame transformation” traditionally means the redefinition of hydrodynamical solutions. In other contexts, mostly a “frame” refers to a coordinate system. In this work, we always distinguish these terminologies by explicitly mentioning “hydrodynamical” in the former contexts. A coordinate transformation also changes the fields when evaluated at the same coordinate value, as the choice of the coordinate system at rest is not unique Kodama:1984ziu . In the present case, a tangent of coordinate transformation is
$\xi^{\mu}=\sum_{n=0}\xi^{\mu}_{(n)},$ (127)
where $n$ counts the order in the covariant derivative. Then, the change of field variables is expressed by the Lie derivative associated with the tangent. The leading order coordinate transformation is, for example,
$x^{\mu}\to x^{\mu}+\pounds_{\xi}x^{\mu}=x^{\mu}+\xi^{\mu}.$ (128)
Note that a coordinate system is a set of 4-scalar fields, $x^{\mu}$. In our case, $\beta$ and $\mu$ are 4-scalar fields, and $u_{\mu}$ is a 1-form, which transform
$\delta\beta$ =£$=\pounds_{\xi}\beta=\xi^{\rho}\nabla_{\rho}\beta,$ (129)
$\delta\mu$ =£$=\pounds_{\xi}\mu=\xi^{\rho}\nabla_{\rho}\mu,$ (130)
$\delta u_{\mu}$ =£$=\pounds_{\xi}u_{\mu}=\xi^{\rho}\nabla_{\rho}u_{\mu}+u_{\rho}\nabla_{\mu}u^{\rho}.$ (131)
In the lowest order in the derivative expansion, we have
$\xi_{(0)}^{\mu}=f(\beta,\mu)u^{\mu}.$ (132)
The “gauge transformation” generated by $\xi_{(0)}^{\mu}$, is
$\delta\beta$ $=fu^{\rho}\nabla_{\rho}\beta,$ (133)
$\delta\mu$ $=-fu^{\rho}\mu\beta^{-1}\nabla_{\rho}\beta+f\beta^{-1}u^{\rho}\nabla_{\rho}(\beta\mu),$ (134)
$\delta u_{\sigma}$ $=fu^{\rho}\nabla_{\rho}u_{\sigma}-(f_{\beta}-f_{\mu}\mu\beta^{-1})\nabla_{\sigma}\beta-f_{\mu}\beta^{-1}\nabla_{\sigma}(\beta\mu),$ (135)
which is a subclass of the field redefinition.
@@ 3 Effective theory for hydrodynamical collective excitations
So far, we have considered hydrodynamical solutions. The path integral for the fast mode is integrated over a given initial state, and $\Psi=\{T,J\}$ is determined. When initial states are thermal, we derived constraints on the generating functional. In this paper, however, we are more interested in the dynamics of collective excitations around hydrodynamical solutions rather than the hydrodynamical solutions themselves. Such collective modes are expressed by the NG modes associated with the symmetry-breaking pattern induced by the external fields. In this section, we consider the effective theory for the NG modes.

### v3

@@ 3.2 Hydrodynamical on-shell transformation
To derive Eqs. (3.29) and (3.30), we have exhausted the freedom of hydrodynamical frame transformations. The transport coefficients with a tilde are hydrodynamical frame invariants, so we cannot further eliminate them by the hydrodynamical frame transformation. However, when evaluated on the lowest-order hydrodynamical solutions, one can further simplify the coefficients [1].
The lowest-order hydrodynamical equations are:
$\nabla_{\mu}(nu^{\mu})$ =0=0,= 0 , (3.38)
$u_{\nu}\nabla_{\rho}(\rho u^{\rho}u^{\nu}+P\gamma^{\rho\nu})$ =0=0,= 0 , (3.39)
$\gamma_{\mu\nu}\nabla_{\rho}(\rho u^{\rho}u^{\nu}+P\gamma^{\rho\nu})+\gamma_{\mu\nu}nu_{\rho}F^{\rho\nu}$ =0=0,= 0 , (3.40)
which can be recast into the following forms:
$u^{\rho}\nabla_{\rho}\beta$ $=\frac{n\rho_{\mu}-n_{\mu}(\rho+P)}{n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu}}\nabla_{\rho}u^{\rho},$ (3.41)
$u^{\rho}\nabla_{\rho}(\beta\mu)$ $=\frac{-\beta n\rho_{\beta}+\mu n\rho_{\mu}+\beta n_{\beta}(\rho+P)-\mu n_{\mu}(\rho+P)}{\beta(n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu})}\nabla_{\rho}u^{\rho},$ (3.42)
$u^{\rho}\nabla_{\rho}u_{\mu}$ $=-\frac{P_{\beta}-\frac{\mu}{\beta}P_{\mu}}{\rho+P}\gamma_{\mu}{}^{\nu}\nabla_{\nu}\beta-\frac{P_{\mu}}{\beta(\rho+P)}\gamma_{\mu}{}^{\nu}\nabla_{\nu}(\beta\mu)-\frac{n}{\rho+P}u^{\rho}F_{\rho\mu}.$ (3.43)
@@
$P_{(1)}$ $={\bm{\zeta}}\nabla_{\rho}u^{\rho},$ (3.44)
$j^{\mu}_{(1)}$ $={\bm{\tau}}\gamma^{\mu\nu}\nabla_{\nu}\beta-{\bm{\chi}}_{1}\beta\gamma^{\mu\nu}\nabla_{\nu}(\beta\mu)-{\bm{\chi}}_{2}\beta^{2}\gamma^{\mu\nu}u^{\rho}F_{\rho\nu},$ (3.45)
${\bm{\zeta}}$ $\equiv\tilde{\zeta}+\left(\frac{\partial P}{\partial\rho}\right)_{n}\tilde{\lambda}+\left(\frac{\partial P}{\partial n}\right)_{\rho}\tilde{\nu},$ (3.46)
${\bm{\tau}}$ $\equiv\tilde{\tau}_{1}-\tilde{\tau}_{2},$ (3.47)
${\bm{\chi}}_{1}$ $\equiv\tilde{\chi}_{1}-\frac{n}{\beta(\rho+P)}\tilde{\tau}_{2},$ (3.48)
${\bm{\chi}}_{2}$ $\equiv\tilde{\chi}_{2}-\frac{n}{\beta(\rho+P)}\tilde{\tau}_{2}.$ (3.49)
Here, we have used Eqs. (B.15) and (B.17). The $\bm{\zeta}$, $\bm{\chi}_{i}$ and $\bm{\tau}_{i}$ are hydrodynamical frame invariants by definition.
When imposing the KMS condition in the original hydrodynamical frame as presented in App. B.2, one can further constrain the coefficients as
${\bm{\tau}}=0,~{\bm{\chi}}_{1}={\bm{\chi}}_{2}\equiv{\bm{\chi}}.$ (3.50)
@@
$x$ =1=1,= 1 , (3.52)
$y$ $=-\left(\frac{\partial P}{\partial\rho}\right)_{n},$ (3.53)
$z$ $=-\left(\frac{\partial P}{\partial n}\right)_{\rho},$ (3.54)
in Eq. (2.70), and
$x$ $=-\frac{n}{\beta(\rho+P)},~y=1,$ (3.55)
in Eq. (2.69). Now, we arrive at the so-called Landau frame, where the first-order transport coefficients are characterized by three independent hydrodynamic parameters: the shear viscosity $\eta$, bulk viscosity $\zeta$, and charge conductivity $\chi$.
In Ref. [23], the ambiguity of the on-shell condition was emphasized. Specifically, it was shown that one can eliminate the term $\nabla_{\mu}u^{\mu}$ and express the constitutive relations in terms of either $u^{\mu}\nabla_{\mu}\beta$ or $u^{\mu}\nabla_{\mu}(\beta\mu)$. These correspond to the conditions $\tilde{\lambda}=\tilde{\zeta}=\tilde{\tau}=0$ and $\tilde{\zeta}=\tilde{\nu}=\tilde{\tau}=0$, respectively, thus modifying the hydrodynamic equations of motion. Therefore, on-shell solutions reintroduce ambiguities in the transport coefficients, which we refer to as on-shell transformations.
@@ 3.3 Traditional hydrodynamical frames
The derivation of the Landau frame presented above differs from the original prescription given by Landau and Lifshitz, which we now discuss.
Suppose one imposes the KMS condition after maximally eliminating coefficients via field redefinition. This corresponds to taking $T_{(1)}^{\mu\nu}u_{\nu}=j_{(1)}^{\mu}u_{\mu}=0$, followed by the imposition of the second law of thermodynamics as outlined in the textbook by Landau and Lifshitz. The first condition leads to $\epsilon=\lambda_{1}=\nu_{1}=\nu_{2}=\nu_{3}=\nu_{4}=\kappa_{1}=\kappa_{2}=\tau_{3}=\tau_{4}=0$. The second condition yields $\lambda_{2}=\lambda_{1}$, $\nu_{5}=\nu_{2}$, $\kappa_{i}=\kappa$, $\tau_{i}=\tau$, and $\chi_{i}=\chi$. As a result, $T^{\mu\nu}$ and $J^{\mu}$ are expressed in terms of three non-vanishing transport coefficients: $\eta$, $\zeta$, and $\chi$. The noise coefficients vanish under the KMS condition except for $w_{4}=\zeta$, $w_{5}=\eta$, and $y_{2}=\chi$. In other words, in the Landau frame, several hydrodynamical invariants are set to zero: $\tilde{\lambda}=\tilde{\nu}=\tilde{\tau}=0$, while $\tilde{\zeta}=\zeta$ and $\tilde{\chi}=\chi$.
Similarly, the Eckart frame is realized by imposing $T_{(1)}^{\mu\nu}u_{\nu}u_{\mu}=j_{(1)}^{\mu}=0$, which leads to $\epsilon=\lambda_{1}=\nu_{1}=\nu_{2}=\nu_{3}=\nu_{4}=\tau_{1}=\tau_{2}=\chi_{1}=\chi_{2}=0$. Then, under the KMS condition, the frame invariants are set to $\tilde{\lambda}=\tilde{\nu}=\tilde{\chi}=0$, with $\tilde{\zeta}=\zeta$ and $\tilde{\tau}=-n\kappa/\beta(\rho+P)$.
These observations make it clear that the Landau and Eckart frames are not equivalent under hydrodynamical transformations, as they yield different hydrodynamical frame invariants. The two frames are related only up to a hydrodynamical frame transformation followed by an on-shell transformation.
At first glance, the freedom to solve lowest-order hydrodynamics arbitrarily might suggest that on-shell transformations are unphysical. However, such transformations alter the nature of the equations of motion and have implications for the analysis of stability and causality [23]. Indeed, earlier studies have concluded that traditional first-order hydrodynamics is neither stable nor causal, largely due to these ambiguities. In Sec. 5, we will demonstrate that the mode analysis remains unaffected by this freedom.
It is also important to note that neither hydrodynamical frame transformations nor on-shell transformations affect the noise sector. Therefore, once the KMS condition is imposed on the most general first-order action, the noise coefficients cannot be further reduced. This apparent inconsistency suggests that a full on-shell treatment of hydrodynamics—including noise fields—is necessary. We will return to this point in Sec. 6, where we show how the noise coefficients can indeed be further constrained.
@@ 3.4 Ambiguity of rest frames
In hydrodynamics, a “frame transformation” traditionally refers to the redefinition of $\beta^{\mu}$. In other contexts, however, the term “frame” typically denotes a coordinate system. In this work, we distinguish these usages by explicitly specifying “hydrodynamical” when referring to the former.
A coordinate transformation changes the field values evaluated at the same coordinate value, since the choice of the coordinate system at rest is not unique [27]. In our setting, the generator of the coordinate transformation can be expanded as
$\xi^{\mu}=\sum_{n=0}\xi^{\mu}_{(n)},$ (3.56)
where $n$ indicates the order in the covariant derivative expansion. The change in field variables is then expressed through the Lie derivative associated with $\xi^{\mu}$. For example, the leading-order coordinate transformation is
$x^{\mu}\to x^{\mu}+\pounds_{\xi}x^{\mu}=x^{\mu}+\xi^{\mu}.$ (3.57)
Note that a coordinate system is represented by a set of scalar fields $x^{\mu}$. In our case, $\beta$ and $\mu$ are scalar fields, and $u_{\mu}$ is a one-form. Their transformations are given by
$\delta\beta$ =£$=\pounds_{\xi}\beta=\xi^{\rho}\nabla_{\rho}\beta,$ (3.58)
$\delta\mu$ =£$=\pounds_{\xi}\mu=\xi^{\rho}\nabla_{\rho}\mu,$ (3.59)
$\delta u_{\mu}$ =£$=\pounds_{\xi}u_{\mu}=\xi^{\rho}\nabla_{\rho}u_{\mu}+u_{\rho}\nabla_{\mu}\xi^{\rho}.$ (3.60)
At the lowest order in the derivative expansion, the transformation generator is given by
$\xi_{(0)}^{\mu}=f(\beta,\mu)u^{\mu}.$ (3.61)
The resulting “gauge transformation” generated by $\xi_{(0)}^{\mu}$ is then
$\delta\beta$ $=f\,u^{\rho}\nabla_{\rho}\beta,$ (3.62)
$\delta\mu$ $=-f\,u^{\rho}\mu\beta^{-1}\nabla_{\rho}\beta+f\,\beta^{-1}u^{\rho}\nabla_{\rho}(\beta\mu),$ (3.63)
$\delta u_{\sigma}$ $=f\,u^{\rho}\nabla_{\rho}u_{\sigma}-(f_{\beta}-f_{\mu}\mu\beta^{-1})\nabla_{\sigma}\beta-f_{\mu}\beta^{-1}\nabla_{\sigma}(\beta\mu),$ (3.64)
which represents a subclass of field redefinitions. In addition, a general diffeomorphism transformation of the metric is given by
$g_{\mu\nu}\to g_{\mu\nu}+\nabla_{\mu}\xi_{\nu}+\nabla_{\nu}\xi_{\mu}.$ (3.65)
Thus, a general coordinate transformation necessarily induces a perturbation of the metric. Although gauge transformations of the metric are unavoidable, this point has not been emphasized in the literature. Therefore, it is important to clearly distinguish general coordinate transformations from field redefinitions of thermodynamical variables.
Section summary.
This section analyzed the twofold redundancy present in relativistic viscous hydrodynamics: the hydrodynamical frame transformation and the on-shell transformation. The former originates from field redefinitions of $\beta^{\mu}$ at first order, which leave the generating functional invariant while altering the form of the constitutive relations. In contrast, the on-shell transformation employs the leading-order equations of motion to further eliminate transport coefficients. Although this reduction does not change physical observables, the physical content must be expressed in terms of the frame-invariant combinations $\eta$, $\bm{\zeta}$, $\bm{\chi}_{1/2}$, and $\bm{\tau}_{1/2}$. After maximally reducing the coefficients using the KMS condition, the number of first-order transport coefficients is reduced to three: $\eta$, $\bm{\zeta}$, $\bm{\chi}$. Nonetheless, it has been pointed out that on-shell transformations can change the character of the equations—from hyperbolic to elliptic—which is a potential source of instability. We also showed that, while these transformations do not affect the noise terms at the order considered, the full noise sector must be consistently solved on-shell to ensure compatibility with the KMS condition. Finally, we clarified the distinction between hydrodynamical and relativistic coordinate frames, showing that leading-order coordinate transformations constitute a subclass of hydrodynamical redefinitions, with a nontrivial transformation on the metric.
@@ 4 Effective theory for hydrodynamical collective excitations
In the previous sections, we constructed the generating functional for a general U(1) symmetric system and constrained its coefficients using the KMS symmetry and the unitarity of the UV theory. We also discussed the redundancies associated with field redefinitions of $\beta^{\mu}$ and on-shell transformations. In this section, we introduce the hydrodynamical mode and construct its effective quadratic action.
A conventional and widely used approach to relativistic hydrodynamics treats $\beta$, $u^{\mu}$, and $\mu$ as thermodynamic variables, and analyzes their linear perturbations under the assumption of local thermal equilibrium. This method has been highly successful in phenomenological applications, as it allows for straightforward analysis based on linearized conservation equations.
However, if one wishes to understand the underlying symmetry structure or systematically derive the dynamics of fluctuations, it is useful to revisit this formulation from the perspective of EFT. In particular, the conventional hydrodynamic perturbations do not transform linearly under the symmetries of the generating functional, and thus the implementation of spacetime covariance—especially under Lorentz boosts—is not manifest.
In the presence of a finite temperature vector $\beta^{\mu}$, which defines a preferred frame, the full Lorentz symmetry is no longer manifest at the level of the background. Nevertheless, the microscopic theory may still be Lorentz invariant, and the preferred frame should be understood as a feature of the ensemble. From this viewpoint, the symmetry breaking is not explicit but rather spontaneous, as the thermal state selects a specific rest frame.
This motivates us to treat the hydrodynamical modes as nonlinear realizations of spontaneously broken spacetime symmetries, similar to Goldstone modes in other EFT contexts. Indeed, sound waves in the fluid can be understood in this way, and their transformation under boosts acquires a nonlinear character. This framework naturally leads to an effective action formulation, which respects the original symmetry structure at the level of the generating functional.
The construction of such an EFT has been initiated in Ref. [15], which introduces the concept of emergent fluid spacetime to realize the full symmetry algebra, including fluctuation and dissipation. While this approach is elegant and powerful, in this work we opt for a simpler formulation that focuses on symmetry realization in physical spacetime, in order to facilitate direct analysis of the hydrodynamical modes themselves.